OpenStudy (anonymous):

you roll a number cube twice. Find P(even, then not 2). Write the probability as a fraction in simplest form.

OpenStudy (anonymous):

dice generally have 6 sides if you role it twice then the are a total of 12 different out comes now what do you mean when you say even then not 2

OpenStudy (kirbykirby):

I think they mean they want to obtain the event : even # on the 1st roll, and then a number other than 2 on the 2nd roll.

OpenStudy (kirbykirby):

P(even) = 3/6 = 1/2 since there are three even numbers: 2, 4, 6 P(not 2) = 5/6, since you can get 1, 3, 4, 5, 6 P(even, not 2) = P(even) * P(not 2) = 1/2 * 5/6 = 5/12